The amount of time a sheep can live after diagnosed with a sickness follows a mean of 3 years and a standard deviation of 1 year. What is the probability it will live more than 6 years?
How many months after would you expect 3.5% of the population to live?
Answer #1

Mu=3, sigma=1

P(X > 6 ) = ?

Z = ( x - mu ) / sigma = ( 6 - 3 ) / 1 = 3

P(z > 3 ) = 1 - P(z<3) = 1 - 0.9987 = 0.0013

P( Z < z ) = 0.035

z = -1.81

x = mu + z * sigma

x = 3 - 1.81 * 1

x = 1.19

x = 1.19 * 12 = 14.28 momths